A closed gas cylinder contains 0.500 mol H2, 0.300 mol O2, 1.200 mol CO2 at a temperature of 25 °C and a pressure of 2.00 atm.

  1. Explain Dalton’s Law of partial pressure.
  2. Calculate the volume of the cylinder.
  3. Calculate the partial pressure of O2 in the container.
1
Expert's answer
2021-04-27T07:50:01-0400

1.According to Dalton's law of partial pressures, the total pressure by a mixture of gases is equal to the sum of the partial pressures of each of the constituent gases



2.Molar Mass of H2 = 2.016

= 0.5 × 2.016 = 1.008g

Molar Mass of O2= 31.996

0.3 × 31.996 = 9.5988g

Molar Mass of CO2=44.01

1.2 × 44.01 = 52.812g

= 1.008 + 9.5988 + 52.812

= 63.4188ml




3.9.5988 × 2 /63.4188

=0.3am